OpenStudy (kkutie7):

Please help me out + tell me what I'm doing wrong. 12a: Two charges are held in place as shown. Determine the magnitude and direction of the electric field at the origin. Use Q = 3 nC, q = −2 nC, D = 4 m, L = 2 m and H = 5 m. b: Determine the magnitude and direction of the acceleration of a 2 µC charge of mass 10−25 kg which is released at the origin.

OpenStudy (kkutie7):



OpenStudy (kkutie7):

For Q: \[E_{Q}=\frac{8.99*10^{9}Nm^{2}}{C^{2}}*\frac{3nC}{(4m)^{2}}*\frac{1*10^{-9}C}{nC}=\frac{1.69N}{C}x\] for q: \[E_{q}=\frac{8.99*10^{9}Nm^{2}}{C^{2}}*\frac{-2nC}{29m^{2}}*\frac{1*10^{-9}C}{nC}=\frac{-0.62N}{C}\] \[R=2mx+5my\] \[r=\frac{R}{|R|}=\frac{2x}{\sqrt{29}}+\frac{5y}{\sqrt{29}}\] \[\frac{-0.62N}{C}(\frac{2}{\sqrt{29}}x+\frac{5}{\sqrt{29}}y)\]

OpenStudy (dan815):

http://prntscr.com/9uf87w

OpenStudy (kkutie7):

the 2 and the 5?

OpenStudy (dan815):

ya or take the 0.62 as positive

OpenStudy (kkutie7):

Oh right let me fix that

OpenStudy (kkutie7):

\[E_{tot}=\frac{1.69N}{C}x+\frac{1.24N}{\sqrt{29}C}x+\frac{3.1N}{\sqrt{29}C}y\] \[E_{tot}=\frac{1.69*\sqrt{29}N+1.24N}{\sqrt{29}C}x+\frac{3.1N}{\sqrt{29}C}y\] \[E_{tot}=\frac{1.920N}{C}x+\frac{0.576N}{C}y\]

OpenStudy (kkutie7):

\[\sqrt{(1.920N/C)^{2}(0.576N/C)^{2}}=2N/C\]

OpenStudy (dan815):

k good (:

OpenStudy (kkutie7):

 \[\theta=tan^{-1}(\frac{0.576}{1.920})=16.7\]

OpenStudy (dan815):

yep gj

OpenStudy (kkutie7):

thank you for the help =)

OpenStudy (dan815):

welcome!

OpenStudy (kkutie7):

b: Determine the magnitude and direction of the acceleration of a 2 µC charge of mass 10^−25kg which is released at the origin.

OpenStudy (dan815):

F=q*E

OpenStudy (kkutie7):

\[F=2\mu C*E=2\mu C*\frac{10^{-6}}{1\mu }*\frac{2N}C{}=4.0*10^{-6}N\]

OpenStudy (kkutie7):

\[F=m*a\] \[\frac{F}{m}=a=\frac{4.0*10^{-6}N}{10^{-25}kg}=4.0*10^{19}\]

OpenStudy (dan815):

yep now apply it to the right direction

OpenStudy (kkutie7):

should be in the same direction as the field no?

OpenStudy (dan815):

yeah

OpenStudy (kkutie7):

sweet. thanks again

OpenStudy (dan815):

have u learnt about complex numbers yet

OpenStudy (dan815):

u dont need to do this, u can just find the x and y lengths for a 16.7 degree angle but a shortcut since u know the angle already \[Acceleration=exp(i*\frac{16.7}{360}*2\pi) ~*~4.0*10^{19}\] the real part of that is in the x component of acc and the imaginary part is the y component of acc just ignore, if it it doesnt make sense. i just think its kind of a neat fast way to get vector representations of magnitudes in any direction quickly once u know the angle u want it to go in

OpenStudy (dan815):

or Acc=4.0*10^19 * cos(16.7deg) x + 4.0*10^19 * sin(16.7deg) y